the product of two consecutive even integers is 168. do not solve.

Found 2 solutions by ewatrrr, jorel1380:
Answer by ewatrrr(24785)

 
Hi,
Let x and (x+2)represent the two consecutive even integers:
Question states***
x(x+2) = 168
fyi
x^2 + 2x -168 = 0
(x+14)(x-12) = 0 x = -14 or x = 12
the two consecutive even integers are 12,14 OR -14,-12

Answer by jorel1380(3719)

n(n+2)=168
n2+2n=168
n2+2n-168=0..